OpenStudy (anonymous):

Will give a medal! Evaluate. 3 – 7 • 5 + 2 • 4 A. -40 B. -24 C. -12 D. 28

OpenStudy (anonymous):

I think its D

OpenStudy (kendricklamar2014):

Don't giver direct answer @Supermariofan153

OpenStudy (kendricklamar2014):

What do you think the answer is @Rockercats622

OpenStudy (anonymous):

?

OpenStudy (kendricklamar2014):

Don't give direct answers means don't say the answer. Show them the steps to get to the answer, but let them say the answer not you.

OpenStudy (anonymous):

oh sorry :\

OpenStudy (idealist10):

3-7*5+2*4=3-35+8=-32+8=-24 Answer: -24, B.